OpenStudy (anonymous):

Solve the system of equations. Write solution as a point (x,y,z). Write no solutioon if inconsistent. 2x-2y+3z=5 2x+y-2z=-1 4x-y-3z=0 (X,y,z) =

OpenStudy (anonymous):

have you done matrices?

OpenStudy (anonymous):

Subtract first from second. Then twice second from third.

OpenStudy (anonymous):

No matrices. So far just elimination and substitution routes.

OpenStudy (anonymous):

okay, then do what estudier suggested

OpenStudy (anonymous):

Subtract the first from tthe second, and then subtract twice the

OpenStudy (anonymous):

yes, subtract the first equation from the second. post what you get.

OpenStudy (anonymous):

3y-5z=-6

OpenStudy (anonymous):

okay good., now subtract twice the first equation from the third equation

OpenStudy (anonymous):

x= 2/3 y=-1/3 z=1

OpenStudy (anonymous):

Great explanation,shadowfax...

OpenStudy (anonymous):

wow shadowfax, you can solve simultaneous equations! Care explaining to ashes2ashes how you did that?

OpenStudy (anonymous):

this forum is not a competition to see if you can solve a problem, it is to explain to the person how to do the problem. There is no achievement in solving a problem you did in high school, there is significant achievement in being able to explain it to a novice.

OpenStudy (anonymous):

next part is -3y+z=2

OpenStudy (anonymous):

then subtract that answer from the first one I found?

OpenStudy (anonymous):

that gives me z=1

OpenStudy (anonymous):

then substitute that back into my equation of 3y-5z=-6

OpenStudy (anonymous):

that leaves me with y= -1/3 correct?

OpenStudy (anonymous):

right. very good now use y and z to find x

OpenStudy (anonymous):

you are helping a ton thanks... l;ets see if I get this

OpenStudy (anonymous):

x= 2/3 then got to 4x=2 2/3 then divide by 4 gives me 2/3

OpenStudy (anonymous):

so (2/3,-1/3,1) I believe

OpenStudy (anonymous):

right. well done!

OpenStudy (anonymous):

thanks! will ahve a similar one to post in as econd to make sure I am doing it right